OpenStudy (anonymous):

solve the equation for x: 4x-8y=16

OpenStudy (anonymous):

add 8y to both sides, and then divide by 4 on both sides

OpenStudy (anonymous):

x = -2y + 4 x = 2y - 4 x = -2y - 4 x = 2y + 4

OpenStudy (anonymous):

4x -8y =16 +8y +8y 4x = 16 + 8y divide by 5 on both sides x = (16+8y)/4 x = 4+2y

OpenStudy (anonymous):

Okay thank you